OpenStudy (anonymous):

Find the arc length of a central angle of 36 degrees in a circle whose radius radius in 2 inches. A.18 in. B. 2 π/5in C.72in. D. π/10in

OpenStudy (michele_laino):

please, first you have to express your angle in radians, namely 36 degrees hwo radians are?

OpenStudy (michele_laino):

oops ...how many radians are?

OpenStudy (michele_laino):

for example 90 degrees are pi/2 radians

OpenStudy (anonymous):

1?

OpenStudy (michele_laino):

no, please if you want to find radians starting from degrees, you have to solve this proportion: \[180:36= \pi :x\] please solve for x

OpenStudy (michele_laino):

do you know how to solve that proportion?

OpenStudy (anonymous):

no

OpenStudy (michele_laino):

please you have to apply the fundamental property of proportions, namely: product of the means equals the product of the extremes, so: 36* pi= 180*x

OpenStudy (michele_laino):

\[36* \pi= 180* x\]

OpenStudy (michele_laino):

now, please divide both sides of that equation by 180

OpenStudy (michele_laino):

namely: \[\frac{ 36* \pi }{ 180 }=\frac{ 180 * x }{ 180 }\] plese simplify that expression

OpenStudy (michele_laino):

oops ...please simplify ...

OpenStudy (anonymous):

36π/180= x?

OpenStudy (michele_laino):

ok! now what is \[\frac{ 180 }{ 36 }=...\]

OpenStudy (anonymous):

5

OpenStudy (michele_laino):

perfect so your angle is: \[x=\frac{ \pi }{ 5 }\]

OpenStudy (michele_laino):

now by definition of measure of an arc, in order to find the length of your arc, please you have to apply this formula: \[arc-length=x * radius\] where x is your angle in radians, so?

OpenStudy (michele_laino):

what is your arc-length, please?

OpenStudy (michele_laino):

\[x* radius= \frac{ \pi }{ 5 }*2=...\]

OpenStudy (anonymous):

B?

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

Thank you!

OpenStudy (michele_laino):

Thank you!

OpenStudy (triciaal):

use the ratio: angle/360 = length/2*pi*r